# Unfitted hybrid high-order methods stabilized by polynomial extension for elliptic interface problems

### v1

@@ Abstract
In this work, we propose the design and the analysis of a novel hybrid high-order (HHO) method on unfitted meshes. HHO methods rely on a pair of unknowns, combining polynomials attached to the mesh faces and the mesh cells. In the unfitted framework, the interface can cut through the mesh cells in a very general fashion, and the polynomial unknowns are doubled in the cut cells and the cut faces. In order to avoid the ill-conditioning issues caused by the presence of small cut cells, the novel approach introduced herein is to use polynomial extensions in the definition of the gradient reconstruction operator. Stability and consistency results are established, leading to optimally decaying error estimates. The theory is illustrated by numerical experiments.
{adjustwidth}
1cm1cm Keywords. Hybrid high-order methods, Unfitted methods, Polynomial extension, Interface problem, Curved boundary. {adjustwidth}1cm1cm Mathematics Subjects Classification. 65N15, 65N30, 65N85.
@@ Introduction
Nevertheless, difficulties remain when using hybrid nonconforming methods in the presence of curved boundaries or interfaces, whereas the treatment of cells with curved faces in dG methods is more straightforward. One possibility is to enrich (with non-polynomial functions) the set of unknowns on curved faces, as in [38] for HHO and in [9] for ncVEM. Another possibility is to work with mesh cells having flat faces, but to build the geometric representation directly into the discretization, in such a way that solution accuracy is not lost although the mesh does not respect the geometry. This idea has been developed first in the framework of unfitted finite element methods. These methods originate from the penalty method [3, 6] and were further developed for conforming FEM in [5]. Combining the ideas of unfitted FEM with Nitsche’s method for interface problems led to a consistent unfitted finite element method for interface problems in [28]. These ideas were then generalized to include much more complex coupling problems including PDEs on surfaces and multiphysics coupling, leading to the cutFEM framework [15]. Unfitted methods have also been considered for dG [7, 31, 34], HDG [23, 35], and HHO [18] methods. A crucial problem for unfitted methods is to handle the instabilities that can occur for certain intersections of the mesh and the interface. In the cutFEM framework using a conforming approximation in each subdomain, this is solved by the addition of certain stabilized terms [13], whereas in dG methods, which are much more flexible with respect to cell geometry, this can be solved by cell agglomeration [36, 31]. This technique can also be used in the context of conforming finite elements [4] (therein called aggregated FEM).
@@
In the present work, we continue the development of unfitted HHO methods. The combination of cell agglomeration and a local weak coupling derived from the HHO framework has been successfully applied to elliptic interface problems [14], incompressible Stokes flows [16], and wave propagation in heterogeneous media [17]. Nevertheless, if these methods have to be integrated in an existing HHO code, the cell-agglomeration procedure becomes a disruptive element since it requires modifying part of the mesh in the interface zone. It is often desirable to keep the underlying mesh fixed and modify instead the algebraic structure of the bulk unknowns. One technique that fits this purpose is the idea of polynomial extension from interior elements to stabilize the method. The idea was first introduced in the context of Lagrange multipliers [29] and then using Nitsche’s method for finite element [33, 19], isogeometric [12] and VEM [11, 10, 30] methods. The purpose of this work is to develop and analyze this technique for hybrid nonconforming methods, focusing on HHO methods. Owing to the close links between HDG, WG, ncVEM and HHO methods, the present results are relevant to the development of these other methods as well. Moreover, we focus on elliptic interface problems, but observe that elliptic problems posed on curved domains can be handled by the same techniques.
@@
Let us briefly outline the main idea. Recall that a central ingredient in HHO methods is a local gradient reconstruction from the local cell and face unknowns. The stability and consistency properties of this reconstructed gradient crucially hinge on a discrete trace inequality on the normal derivatives at the cell faces. In the case of ill-cut cells, the constant in this inequality degenerates. One remedy, explored in [18, 14], is to use cell agglomeration around ill-cut cells. This procedure produces an aggregated mesh solely composed of well-cut cells, whose size is close to that of the original mesh. The remedy explored herein is to use the cell and face polynomials of ill-cut cells in the gradient reconstruction of a neighboring well-cut cell. With this approach, the mesh cells remain unmodified, but the stencil associated with the gradient reconstruction operator is slightly extended. The main question is then whether the reconstructed gradient maintains the above stability and optimal approximation properties. The proof of these properties is by no means straightforward and is the main contribution of this work. Interestingly, the main analysis tools (discrete trace inequality and polynomial approximation) are of independent interest.
@@
Our goal is to approximate the solution $u\in H^{1}(\Omega_{1}\cup\Omega_{2})$ of the following elliptic interface problem:
[…]
@@
$\text{Find }u\in V_{g_{D}},\qquad a(u,w)=\ell(w)\qquad\forall w\in V_{0},$ (2)
@@
Let $\mathcal{T}_{h}$ be a mesh that covers $\Omega$ exactly. A generic mesh cell is denoted $T\in\mathcal{T}_{h}$, it is considered to be an open set, $h_{T}$ denotes its diameter, and $\boldsymbol{n}_{T}$ its unit outward normal. The mesh size is defined as $h:=\max_{T\in\mathcal{T}_{h}}h_{T}$. For all $T\in\mathcal{T}_{h}$, its neighboring layers $\Delta_{j}(T)\subset\mathcal{T}_{h}$ are defined by induction as $\Delta_{0}(T):=T$ and $\Delta_{j}(T):=\left\{T^{\prime}\in\mathcal{T}_{h}\mid\overline{T^{\prime}}\cap\overline{\Delta_{j-1}(T)}\neq\emptyset\right\}$ for all $j\in\mathcal{N}$. In what follows, we use $j\in\{1,2\}$. To alleviate the notation, we set $\Delta(T):=\Delta_{1}(T)$. We make the mild assumption that, for every mesh cell $T\in\mathcal{T}_{h}$, its convex hull, $\operatorname*{conv}(T)$, satisfies
@@
The mesh $\mathcal{T}_{h}$ is typically composed of cells having a simple shape (although this is not a necessary assumption); in our illustrations, we consider a structured mesh composed of squares. Quite importantly, the mesh $\mathcal{T}_{h}$ is not fitted to the interface $\Gamma$, so that $\Gamma$ can cut arbitrarily across any mesh cell. For all $T\in\mathcal{T}_{h}$, we set $T^{\Gamma}:=T\cap\Gamma$ and $T^{i}:=T\cap\Omega^{i}$ for all $i\in\{1,2\}$ (some of these sets may be empty). The first obvious partition among the mesh cells is between uncut and cut cells. Moreover, among cut cells, we distinguish between well-cut and ill-cut cells. Specifically, we fix a parameter $\vartheta\in(0,1)$ and say that a cut cell $T$ is a well-cut cell if, for all $i\in\{1,2\}$, $T^{i}$ contains a ball of radius $\vartheta h_{T}$, whereas $T$ is an ill-cut cell if this condition fails. It is shown in [18, Lemma 6.2] that, it the mesh size $h$ is small enough (with respect to the curvature of the interface) and the parameter $\vartheta$ is small enough (with respect to the shape-regularity parameter of the mesh), then, for every ill-cut cell, the above condition fails on one and only one side of $T$. In what follows, we assume that this is always the case. The above considerations lead to the following partitions:
@@
[…]
@@
The polynomial extension technique we propose relies on a pairing operator, which maps every ill-cut cell $T\in\mathcal{T}_{h}^{\textup{KO}}$ to an uncut or a well-cut cell that belongs to the neighborhood $\Delta(T)$. We denote the pairing operator as $\mathcal{N}:\mathcal{T}_{h}^{\textup{KO}}\longrightarrow\mathcal{T}_{h}$, so that we have
@@
The pairing operator is essentially the one constructed in [14, Algorithm 1]. One difference, however, is that after Step 3 of this algorithm, we need to add an additional Step 4 performing the same operations as in Step 3, but for any mesh cell in $\mathcal{T}_{h}^{\textup{KO},2}$. One way of proceeding is to set $\mathcal{N}_{2}(T):=S$, whenever $T:=\mathcal{N}_{1}(S)\in\mathcal{T}_{h}^{\textup{KO},2}$ for some cell $S\in\mathcal{T}_{h}^{\textup{KO},1}$. Thus, the well-cut part of $S$ in $\Omega_{2}$ will be used for $T$, and the well-cut part of $T$ in $\Omega_{1}$ will be used for $S$. The reason why Step 4 is not needed in the context of cell-agglomeration is that by agglomerating $T$ and $S$ in Step 3, $T$ does not require any further stabilization. An illustration of the pairing operators $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$ is shown in Figure 3, and two concrete examples are displayed in the central column of Figure 5 below. Notice that it may happen that some cells in $\mathcal{T}_{h}^{\textup{OK}}$ are in the image of both $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$.
@@
The mesh faces are collected in the set $\mathcal{F}_{h}$, and we set $\mathcal{F}_{h}^{i}:=\{F\in\mathcal{F}_{h}\mid F\subset\Omega_{i}\}$ for all $i\in\{1,2\}$. For every mesh face $F\in\mathcal{F}_{h}^{\Gamma}:=\mathcal{F}_{h}\setminus(\mathcal{F}_{h}^{1}\cup\mathcal{F}_{h}^{2})$ cut by the interface $\Gamma$, we define its two sub-faces as $F^{i}:=F\cap\Omega_{i}$ for all $i\in\{1,2\}$. For all $T\in\mathcal{T}_{h}$, $\mathcal{F}_{\partial T}$ is the collection of mesh faces composing the boundary $\partial T$. Moreover, for every cut cell $T\in\mathcal{T}_{h}^{\textup{OK}}\cup\mathcal{T}_{h}^{\textup{KO}}$, recalling that $T^{\Gamma}=T\cap\Gamma$, the boundary $\partial(T^{i})$ of the sub-cell $T^{i}$ is decomposed as
@@
HHO methods utilize polynomial unknowns attached to the mesh cells and to the mesh faces. We consider here the mixed-order HHO method, where the degree of the face unknowns is $k\geq 0$ and that of the cell unknowns is $(k+1)$. For the motivation of using a mixed-order setting, we refer the reader to [18, Remark 4.1].
@@
Here, $\mathbb{P}^{k+1}(T;\mathbb{R})$ (resp., $\mathbb{P}^{k}(F;\mathbb{R})$) consists of the restriction to $T$ (resp., $F$) of scalar-valued $d$ variate polynomials of degree at most $(k+1)$ (resp., $(d-1)$-variate polynomials of degree at most $k$ composed with any affine geometric mapping from the hyperplane supporting $F$ to $\mathbb{R}^{d-1}$). For every cut cell $T\in\mathcal{T}_{h}^{\textup{OK}}\cup\mathcal{T}_{h}^{\textup{KO}}$, following [18, 14], we double the HHO unknowns so as to have the usual unknowns available in each sub-cell, except on $T^{\Gamma}$ where there are no unknowns. Thus, the local HHO unknowns in every cut cell are
@@
with $\llbracket u_{T}\rrbracket_{\Gamma}:=u_{T^{1}}|_{T^{\Gamma}}-u_{T^{2}}|_{T^{\Gamma}}$ and $\delta_{i1}$ denotes the Kronecker delta. Notice that the two gradients are reconstructed independently, but $\boldsymbol{G}_{T^{1}}^{k}$ depends on the cell components on both sides of the interface owing to the last term in (17). This term is only present when $i=1$ because we assumed $\kappa_{1}\leq\kappa_{2}$ (to fix the ideas, otherwise the term is present only when $i=2$). As discussed in [14, Section 2.5], this is instrumental to achieve robustness of the error estimate in the highly contrasted case $\kappa_{1}\ll\kappa_{2}$.
@@
We can now define, for all $T\in\mathcal{T}_{h}$, the local gradient reconstruction operators $\boldsymbol{G}_{T^{i}}^{k}:\widehat{\mathcal{U}}_{T}^{\mathcal{N}}\rightarrow\mathbb{P}^{k}(T^{i};\mathbb{R}^{d})$, for all $i\in\{1,2\}$. For a polynomial $\boldsymbol{q}\in\mathbb{P}^{k}(T^{i};\mathbb{R}^{d})$ with $i\in\{1,2\}$, $\boldsymbol{q}^{+}$ denotes its extension to $T^{i}\cup\bigcup_{S\in\mathcal{N}_{i}^{-1}(T)}S^{i}$. Then, for all $\hat{u}_{T}^{\mathcal{N}}\in\widehat{\mathcal{U}}_{T}^{\mathcal{N}}$,
@@
$\bf\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\textbf{S}^{\Gamma}$$\tiny\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}\boldsymbol{S^{\bar{\imath}}}$$\tiny\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}\boldsymbol{S^{i}}$$\tiny\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\boldsymbol{T^{i}}$$\tiny\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}\boldsymbol{\mathcal{N}_{i}(S)^{i}}$
@@
$\bf\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S_{2}^{\Gamma}$$\bf\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}T^{\Gamma}$$\bf\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S_{1}^{\Gamma}$$\tiny\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\boldsymbol{S^{\bar{\imath}}_{1}}$$\tiny\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}\boldsymbol{S^{i}_{1}}$$\tiny\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}\boldsymbol{T^{i}}$$\tiny\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\boldsymbol{T^{\bar{\imath}}}$$\tiny\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}\boldsymbol{\mathcal{N}_{i}(S_{1})^{i}}$$\tiny\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}\boldsymbol{\mathcal{N}_{\bar{\imath}}(S_{2})^{\bar{\imath}}}$$\tiny\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\boldsymbol{S^{\bar{\imath}}_{2}}$$\tiny\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}\boldsymbol{S^{i}_{2}}$
@@
$\bf\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}T^{\Gamma}$$\bf\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S^{\Gamma}$$\tiny\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\boldsymbol{T^{\bar{\imath}}}$$\tiny\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}\boldsymbol{T^{i}}$$\tiny\color[rgb]{.25,.25,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.25,.25,.25}\pgfsys@color@gray@stroke{.25}\pgfsys@color@gray@fill{.25}\boldsymbol{S^{i}}$$\tiny\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\boldsymbol{S^{\bar{\imath}}}$$\tiny\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}\boldsymbol{\mathcal{N}_{i}(T)^{i}}$$\tiny\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}\boldsymbol{\mathcal{N}_{\bar{\imath}}(S)^{\bar{\imath}}}$
@@
$\mathcal{P}_{h}^{\textup{OK}}$ :=$:=\{(T,\iota(T))\,\mid\,T\in\mathcal{T}_{h}^{\textup{uncut}}\}\cup\{(T,1),(T,2)\,\mid\,T\in\mathcal{T}_{h}^{\textup{OK}}\}\cup\{(T,\bar{\imath}),\,\mid\,T\in\mathcal{T}_{h}^{\textup{KO}},i:=\iota(T)\},$ (22a)
$\mathcal{P}_{h}^{\textup{KO}}$ :=$:=\{(T,i)\,\mid\,T\in\mathcal{T}_{h}^{\textup{KO}},i:=\iota(T)\}.$ (22b)
@@
The third stabilization bilinear form is specific to the present setting using polynomial extensions. Its aim is to provide some control on the cell components in ill-cut cells. Specifically, we set, for all $\hat{v}_{h},\hat{w}_{h}\in\widehat{\mathcal{U}}_{h0}$,
$s_{h}^{\mathcal{N}}(\hat{v}_{h},\hat{w}_{h}):=\sum_{(T,i)\in\mathcal{P}_{h}^{\textup{OK}}}\sum_{S\in\mathcal{N}_{i}^{-1}(T)}\kappa_{i}h_{S}^{-1}(v_{S^{i}}-v_{T^{i}}^{+},w_{S^{i}}-w_{T^{i}}^{+})_{(\partial S)^{i}},$ (30)
where the superscript + denotes the extension of a polynomial originally defined on $T^{i}$ to $T^{i}\cup\bigcup_{S\in\mathcal{N}_{i}^{-1}(T)}S^{i}$.
@@
To prepare for the consistency error analysis, we state a useful result on the discrete bilinear and linear forms.

### v2

@@ Abstract
In this work, we study the design and analysis of a novel hybrid high-order (HHO) method on unfitted meshes. HHO methods rely on a pair of unknowns, combining polynomials attached to the mesh faces and the mesh cells. In the unfitted framework, the interface can cut through the mesh cells in a very general fashion, and the polynomial unknowns are doubled in the cut cells and the cut faces. In order to avoid the ill-conditioning issues caused by the presence of small cut cells, the novel approach introduced herein is to use polynomial extensions in the definition of the gradient reconstruction operator. Stability and consistency results are established, leading to optimally decaying error estimates. The theory is illustrated by numerical experiments.
keywords:
Hybrid high-order methods, Unfitted methods, Polynomial extension, Interface problem, Curved boundary.
@@ Introduction
Nevertheless, difficulties remain when using hybrid nonconforming methods in the presence of curved boundaries or interfaces, whereas the treatment of cells with curved faces in dG methods is more straightforward. One possibility is to enrich (with non-polynomial functions) the set of unknowns on curved faces, as in [35] for HHO and in [9] for ncVEM. Another possibility is to work with mesh cells having flat faces, but to build the geometric representation directly into the discretization, in such a way that solution accuracy is not lost although the mesh does not respect the geometry. This idea has been developed first in the framework of unfitted finite element methods. These methods originate from the penalty method [3, 6] and were further developed for conforming FEM in [5]. Combining the ideas of unfitted FEM with Nitsche’s method for interface problems led to a consistent unfitted finite element method for interface problems in [25]. These ideas were then generalized to include more complex coupling problems including PDEs on surfaces and multiphysics coupling, leading to the cutFEM framework [13]. Unfitted methods have also been considered for dG [7, 27, 30], HDG [20, 32], and HHO [16] methods. A crucial problem for unfitted methods is to handle the instabilities that can occur for certain intersections of the mesh and the interface. In the cutFEM framework using a conforming approximation in each subdomain, this is solved by the addition of certain stabilized terms [11], whereas in dG methods, which are more flexible with respect to cell geometry, this can be solved by cell agglomeration [33, 27]. This technique can also be used in the context of conforming finite elements [4] (therein called aggregated FEM).
@@
In the present work, we continue the development of unfitted HHO methods. The combination of cell agglomeration and a local weak coupling derived from the HHO framework has been successfully applied to elliptic interface problems [12], incompressible Stokes flows [14], and wave propagation in heterogeneous media [15]. Nevertheless, if these methods have to be integrated in an existing HHO code, the cell-agglomeration procedure becomes a disruptive element since it requires modifying part of the mesh in the interface zone. It is often desirable to keep the underlying mesh fixed and modify instead the algebraic structure of the bulk unknowns. One technique that fits this purpose is the idea of polynomial extension from interior elements to stabilize the method. The idea was first introduced in the context of Lagrange multipliers [26] and then using isogeometric approximation and Nitsche’s method in [10]. The purpose of this work is to develop and analyze this technique for hybrid nonconforming methods, focusing on HHO methods. Owing to the close links between HDG, WG, ncVEM and HHO methods, the present results are relevant to the development of these other methods as well. Moreover, we focus on elliptic interface problems, but observe that elliptic problems posed on curved domains can be handled by the same techniques.
@@
Let us briefly outline the main idea. Recall that a central ingredient in HHO methods is a local gradient reconstruction from the local cell and face unknowns. The stability and consistency properties of this reconstructed gradient crucially hinge on a discrete trace inequality on the normal derivatives at the cell faces. In the case of ill-cut cells, the constant in this inequality degenerates. One remedy, explored in [16, 12], is to use cell agglomeration around ill-cut cells. This procedure produces an aggregated mesh solely composed of well-cut cells, whose size is close to that of the original mesh. The remedy explored herein is to use the cell and face polynomials of ill-cut cells in the gradient reconstruction of a neighboring well-cut cell. With this approach, the mesh cells remain unmodified, but the stencil associated with the gradient reconstruction operator is slightly extended. Moreover, the cell polynomials of the ill-cut cells are stabilized by using the cell polynomials of the associated well-cut cells, in the same spirit as the so-called direct ghost penalty method devised in [31, 28]. The main question is then whether the resulting method maintains the above stability and optimal approximation properties. The proof of these properties is by no means straightforward and is the main contribution of this work. Interestingly, the main analysis tools (discrete trace inequality and polynomial approximation) are of independent interest.
@@
Our goal is to approximate the exact solution ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u^{\textup{ex}}}\in H^{1}(\Omega_{1}\cup\Omega_{2})$ of the following elliptic interface problem:
$-\nabla{\cdot}(\kappa\nabla{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u^{\textup{ex}}})$ $=f$ $\text{ in }\Omega_{1}\cup\Omega_{2},$ (1a)
$\llbracket{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u^{\textup{ex}}}\rrbracket_{\Gamma}$ $=g_{D}$ $\text{ on }\Gamma,$ (1b)
$\llbracket\kappa\nabla{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u^{\textup{ex}}}\rrbracket_{\Gamma}\cdot\boldsymbol{n}_{\Gamma}$ $=g_{N}$ $\text{ on }\Gamma,$ (1c)
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u^{\textup{ex}}}$ $=0$ $\text{ on }\partial\Omega,$ (1d)
@@
$\text{Find }{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u^{\textup{ex}}}\in V_{g_{D}},\qquad a({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u^{\textup{ex}}},w)=\ell(w)\qquad\forall w\in V_{0},$ (2)
@@
[…]
@@
The mesh $\mathcal{T}_{h}$ is typically composed of cells having a simple shape (although this is not a necessary assumption); in our illustrations, we consider a structured mesh composed of squares. Quite importantly, the mesh $\mathcal{T}_{h}$ is not fitted to the interface $\Gamma$, so that $\Gamma$ can cut arbitrarily across any mesh cell. For all $T\in\mathcal{T}_{h}$, we set $T^{\Gamma}:=T\cap\Gamma$ and $T^{i}:=T\cap\Omega^{i}$ for all $i\in\{1,2\}$ (some of these sets may be empty). The current implementation assumes that both sub-cells $T^{i}$ are connected, although this assumption is not needed from a theoretical viewpoint.
The first obvious partition among the mesh cells is between uncut and cut cells. Moreover, among cut cells, we distinguish between well-cut and ill-cut cells. Specifically, we fix a parameter $\vartheta\in(0,1)$ and say that a cut cell $T$ is a well-cut cell if, for all $i\in\{1,2\}$, $T^{i}$ contains a ball of radius $\vartheta h_{T}$, whereas $T$ is an ill-cut cell if this condition fails. It is shown in [16, Lemma 6.2] that, it the mesh size $h$ is small enough (with respect to the curvature of the interface) and the parameter $\vartheta$ is small enough (with respect to the shape-regularity parameter of the mesh), then, for every ill-cut cell, the above condition can fail on one and only one side of $T$. In what follows, we assume that this is always the case. The above considerations lead to the following partitions:
@@
$\mathcal{T}_{h}^{1}$$\mathcal{T}_{h}^{2}$$\mathcal{T}_{h}^{\textup{KO},1}$$\mathcal{T}_{h}^{\textup{KO},2}$$\mathcal{T}_{h}^{\textup{OK}}$$\boldsymbol{\textbf{T}^{1}}$$\boldsymbol{\textbf{T}^{2}}$$(\partial\text{T})^{1}$$(\partial\text{T})^{2}$$\boldsymbol{\textbf{T}^{\Gamma}}$
@@
The polynomial extension technique we propose relies on a pairing operator, which maps every ill-cut cell $S\in\mathcal{T}_{h}^{\textup{KO}}$ to an uncut or a well-cut cell that belongs to the neighborhood $\Delta(S)$. We denote the pairing operator as $\mathcal{N}:\mathcal{T}_{h}^{\textup{KO}}\longrightarrow\mathcal{T}_{h}$, so that we have
@@
The pairing operator is essentially the one constructed in [12, Algorithm 1]. One difference, however, is that after Step 3 of this algorithm, we need to add an additional Step 4 performing the same operations as in Step 3, but for any mesh cell in $\mathcal{T}_{h}^{\textup{KO},2}$. One way of proceeding is to set $\mathcal{N}_{2}(T):=S$, whenever $T:=\mathcal{N}_{1}(S)\in\mathcal{T}_{h}^{\textup{KO},2}$ for some cell $S\in\mathcal{T}_{h}^{\textup{KO},1}$. Thus, the well-cut part of $S$ in $\Omega_{2}$ will be used for $T$, and the well-cut part of $T$ in $\Omega_{1}$ will be used for $S$. The reason why Step 4 is not needed in the context of cell-agglomeration is that by agglomerating $T$ and $S$ in Step 3, $T$ does not require any further stabilization. An illustration of the pairing operators $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$ is shown in Figure 3, and two concrete examples are displayed in the central column of Figure 5 below. Notice that it may happen that some cells in $\mathcal{T}_{h}^{\textup{OK}}$ are in the image of both $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$. The above choice in Step 4 aims at reducing the possible impact of the pairing operator on the resulting stiffness matrix. There may be room for some further improvement, but we do not delve on this aspect here since the impact of the pairing operator on the stencil is already quite modest, as highlighted in Figure 8 below.
@@
The mesh faces are collected in the set $\mathcal{F}_{h}$, and we set $\mathcal{F}_{h}^{i}:=\{F\in\mathcal{F}_{h}\mid F\subset\Omega_{i}\}$ for all $i\in\{1,2\}$. For every mesh face $F\in\mathcal{F}_{h}^{\Gamma}:=\mathcal{F}_{h}\setminus(\mathcal{F}_{h}^{1}\cup\mathcal{F}_{h}^{2})$ cut by the interface $\Gamma$, we define its two sub-faces as $F^{i}:=F\cap\Omega_{i}$ for all $i\in\{1,2\}$. For all $\in\mathcal{T}_{h}, T\mathcal{F}_{\partial T}$ is the collection of mesh faces composing the boundary $\partial T$. Moreover, for every cut cell $T\in\mathcal{T}_{h}^{\textup{OK}}\cup\mathcal{T}_{h}^{\textup{KO}}$, recalling that $T^{\Gamma}=T\cap\Gamma$, the boundary $\partial(T^{i})$ of the sub-cell $T^{i}$ is decomposed as
@@
HHO methods utilize polynomial unknowns attached to the mesh cells and to the mesh faces. We consider here the mixed-order HHO method, where the degree of the face unknowns is $k\geq 0$ and that of the cell unknowns is $(k+1)$. For the motivation of using a mixed-order setting, we refer the reader to [16, Remark 4.1].
@@
Here, $\mathbb{P}^{k+1}(T;\mathbb{R})$ (resp., $\mathbb{P}^{k}(F;\mathbb{R})$) consists of the restriction to $T$ (resp., $F$) of scalar-valued $d$ variate polynomials of degree at most $(k+1)$ (resp., (d1)(d-1)-variate polynomials of degree at most $k$ composed with any affine geometric mapping from the hyperplane supporting $F$ to $\mathbb{R}^{d-1}$). For every cut cell $T\in\mathcal{T}_{h}^{\textup{OK}}\cup\mathcal{T}_{h}^{\textup{KO}}$, following [16, 12], we double the HHO unknowns so as to have the usual unknowns available in each sub-cell, except on $T^{\Gamma}$ where there are no unknowns. Thus, the local HHO unknowns in every cut cell are
@@
with $\llbracket u_{T}\rrbracket_{\Gamma}:=u_{T^{1}}|_{T^{\Gamma}}-u_{T^{2}}|_{T^{\Gamma}}$ and $\delta_{i1}$ denotes the Kronecker delta. Notice that the two gradients are reconstructed independently, but $\boldsymbol{G}_{T^{1}}^{k}$ depends on the cell components on both sides of the interface owing to the last term in (17). This term is only present when $i=1$ because we assumed $\kappa_{1}\leq\kappa_{2}$ (to fix the ideas, otherwise the term is present only when $i=2$). As discussed in [12, Section 2.5], this is instrumental to achieve robustness of the error estimate in the highly contrasted case $\kappa_{1}\ll\kappa_{2}$.
@@
We can now define, for all $T\in\mathcal{T}_{h}$, the local gradient reconstruction operators $\boldsymbol{G}_{T^{i}}^{k}:\widehat{\mathcal{U}}_{T}^{\mathcal{N}}\rightarrow\mathbb{P}^{k}(T^{i};\mathbb{R}^{d})$, for all $i\in\{1,2\}$. For all $(T,i)\in\mathcal{P}_{h}$ and every polynomial $\boldsymbol{q}\in\mathbb{P}^{k}(T^{i};\mathbb{R}^{d})$, $\boldsymbol{q}^{+}$ denotes its extension to $\Delta(T)$ (observe that $T^{i}\cup\bigcup_{S\in\mathcal{N}_{i}^{-1}(T)}S^{i}\subset\Delta(T)$). Then, for all $\hat{u}_{T}^{\mathcal{N}}\in\widehat{\mathcal{U}}_{T}^{\mathcal{N}}$,
@@
[…]
@@
[…]
@@
[…]
@@
$\mathcal{P}_{h}^{\textup{OK}}$ $:=\{(T,\iota(T))\,\mid\,T\in\mathcal{T}_{h}^{\textup{uncut}}\}\cup\{(T,1),(T,2)\,\mid\,T\in\mathcal{T}_{h}^{\textup{OK}}\}\cup\{(T,\bar{\imath}),\,\mid\,T\in\mathcal{T}_{h}^{\textup{KO}},i:=\iota(T)\},$ (22a)
$\mathcal{P}_{h}^{\textup{KO}}$ $:=\{(T,i)\,\mid\,T\in\mathcal{T}_{h}^{\textup{KO}},i:=\iota(T)\}.$ (22b)
@@
The third stabilization bilinear form is specific to the present setting using polynomial extensions. Its aim is to provide some control on the cell components in ill-cut cells and is devised in the same spirit as the so-called direct ghost penalty method from [31, 28]. Specifically, we set, for all $\hat{v}_{h},\hat{w}_{h}\in\widehat{\mathcal{U}}_{h0}$,
$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}s_{h}^{\mathcal{N}}(\hat{v}_{h},\hat{w}_{h}):=\sum_{(T,i)\in\mathcal{P}_{h}^{\textup{OK}}}\sum_{S\in\mathcal{N}_{i}^{-1}(T)}\eta_{\mathcal{N}}\kappa_{i}h_{T}^{-2}(v_{S^{i}}-v_{T^{i}}^{+},w_{S^{i}}-w_{T^{i}}^{+})_{T^{i}},$} (30)
recalling that the superscript + denotes the extension of a polynomial originally defined on $T^{i}$ to $\Delta(T)$, and $\eta_{\mathcal{N}}>0$ is a user-defined parameter.
@@
To prepare for the consistency error analysis, we state a useful result on the discrete bilinear and linear forms. Although this lemma is only used in Section 4, it is given here so as to highlight why the definitions (24)-(25) of the reconstructed gradient are meaningful.